# Robustness and Cybersecurity in the EU Artificial Intelligence Act

### v1

@@ Abstract.
copyright: nonejournalyear: 2025
@@ 1. Introduction
The European Union (EU) recently adopted the Artificial Intelligence Act (AIA) [^1] EU Regulation 2024/1689, 12.7.2025. which creates a legal framework for the development, deployment, and use of “human-centered and trustworthy artificial intelligence (AI)” (Art. 1 AIA). The AIA outlines desirable “ethical principles” of AI systems (Rec. (27)) and, i.a., imposes some of these as legally binding requirements for high-risk AI systems (HRAIS), e.g, AI systems intended to be used to take university admission decisions, or to evaluate individuals’ creditworthiness, and for general-purpose AI models (GPAIMs), e.g., multimodal large language models. While the AIA is recognized as being one of the first legally binding regulatory frameworks for AI (Chee and Hummel, 2024), it has faced criticism for its imprecise and incoherent terminology (Laux et al., 2024; Bomhard and Siglmüller, 2024), which will complicate its practical implementation. Previous work has examined the AIA and its legislative history to clarify terms like explainability (Bordt et al., 2022; Vitali, 2022; Pavlidis, 2024) and fairness (Deck et al., 2024). So far, little attention has been paid to robustness and cybersecurity. Only AI systems classified as high-risk (HRAIS) must meet the robustness and cybersecurity requirements set out in Art. 15 AIA. This paper thus focuses on requirements for HRAIS. To provide a clearer understanding, we compare these requirements with requirements for specific AI models, namely for GPAIMs with systemic risk, in Art. 55 AIA.
@@
Technical solutions to ensure the robustness and cybersecurity of AI systems are often developed within the ML domain. Therefore, it is essential to inform ML research about the legal requirements to ensure compliance with the AIA. However, the vagueness of requirements for cybersecurity and robustness under the AIA makes it challenging to inform ML practitioners about the specific legal requirements to further the development of solutions that can ensure compliance with the AIA. A common understanding between technical and legal domains can be facilitated through technical standards. While the AIA sets out general rules, technical standards specify these rules in detail. Standards are technical specifications designed to provide voluntary technical or quality specifications for current or future products, processes or services. [^2] Art. 1, 2(1) EU Regulation 1025/2012, OJ L 316, 14.11.2012. They prescribe technical requirements, including characteristics such as quality or performance levels, terminology, and test methods. [^3] Art. 2(4)(a) and (c) ibid. Standards have long been integral to EU product legislation under the New Legislative Framework, upon which the AIA is built (Gorywoda, 2009). If approved by the EU Commission, technical standards become harmonized technical standards, which grants a presumption of conformity to products or processes that adhere to them. Consequently, compliance with these standards is deemed to fulfill the requirements of the AIA, thereby incentivizing providers to adopt them (Art. 40 AIA). The development of harmonized technical standards for the AIA has been initiated by the EU Commission and is expected to be completed in the next few years. In addition, the EU Commission is also tasked with developing additional guidelines on the practical implementation on the application of Art. 15 AIA (Art. 96(1)(a) AIA).
@@
  • We analyze and explain the legal requirements related to robustness and cybersecurity in the AIA, identify related shortcomings, and offer possible solutions for some of these shortcomings.
@@
This paper is structured as follows: Section 2 provides a short background on robustness and cybersecurity in the ML literature. Section 3 provides an introduction to the AIA and Art. 15 AIA. Section 5 analyzes the requirements outlined in Art. 15 AIA for HRAIS, addressing both general challenges pertinent to robustness and cybersecurity, as well as specific issues related to each requirement. Section 6 examines the requirements in Art. 55 AIA relevant to GPAIMs with systemic risk. Section 7 concludes with a summary and recommendations for future research.
@@ 2. An ML Perspective on Robustness and Cybersecurity
From a technical standpoint, adversarial robustness is one aspect of cybersecurity. Research on cybersecurity focuses on developing defenses that protect computer systems from attacks compromising their confidentiality, integrity, or availability (Dasgupta et al., 2022). This encompasses aspects like data storage, information access and modification, and secure data transmission over networks (Sarker et al., 2021). Unlike robustness, cybersecurity is not a stand-alone concept in ML, but is discussed more broadly as both a tool for ensuring cybersecurity and a potential source of cybersecurity risks. [^4] E.g., NeurIPS’18 Workshop on Security in ML (Papernot et al., 2018), ICML’22 Workshop on ML for Cybersecurity (Applebaum et al., 2022). ML algorithms can be employed to detect and mitigate cybersecurity threats (Sarker et al., 2021), but can also introduce specific vulnerabilities that adversaries may exploit, such as data poisoning or adversarial attacks (Roshanaei et al., 2024; Rosenberg et al., 2021).
@@
The AIA creates harmonized rules for certain AI systems in order to incentivize the use of such systems in the EU market and prevent regulatory fragmentation between member states (Rec. (1)). It is formally structured into recitals (Rec.), articles (Art.), and annexes. Recitals are not strictly legally binding and outline the rationale behind the articles, articles delineate specific binding obligations, and the annexes provide additional details and specifications to support the articles  (Klimas and Vaiciukaite, 2008). [^5] From now on, whenever we cite recitals, we refer to those in the AIA without explicitly indicating it. Art. 3(1) AIA defines an AI system as “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs […] that can influence physical or virtual environments”. These AI systems are regulated differently based on their perceived risk level (Bomhard and Siglmüller, 2024; Sioli, 2021): Those posing unacceptable risks, such as social scoring, are prohibited or subject to qualified prohibitions; HRAIS, such as those used in medical devices, are allowed but must comply with certain requirements and undergo pre-assessment; other AI systems are subject only to specific transparency and information obligations. Among these categories, only HRAIS must fulfill the robustness and cybersecurity requirements under Art. 15 AIA. According to Art. 16(a) AIA, providers of HRAIS must ensure compliance with these requirements. An AI system is considered a HRAIS if it is either a safety component of a product or a product itself regulated under specific legislation, such as medical devices, machinery, or toys (Art. 6(1) AIA, Annex I), or if it poses a significant risk of harm to the health, safety, or fundamental rights of individuals in specific areas, such as education, employment, or law enforcement (Art. 6(2) and (3) AIA, Annex III).
@@
In addition to AI systems, the AIA establishes a separate regime of legal requirements in chapter V of the AIA for a very specific type of AI models, namely GPAIM (e.g., multimodal large language models, see also Section 6). GPAI models are AI models that can perform tasks that they were not originally trained for (Gutierrez et al., 2023), such as large language models (OpenAI, 2023; Gemini Team et al., 2023), or large text-to-image models (Ramesh et al., 2022). Other types of AI models are not regulated and mentioned in the AIA. These are models that are created with a specific objective and can only accomplish tasks they are trained to perform (e.g., translation, classification).
@@
The AIA does not directly define specific technical requirements. Instead, it sets out ’essential requirements’ that AI systems must comply with, and which are concretized by so-called technical standards. The regulatory concept of relying on standardization is a well-established process in EU product legislation and is referred to as the New Legislative Framework (Gorywoda, 2009). It traditionally involves the participation of stakeholders, such as providers of AI systems. The European Commission has already issued a standardization request to the European Committee for Standardisation and the European Committee for Electrotechnical Standardisation to develop standards for the AIA until 30 April 2025.666C(2023)3215 - Standardisation request M/593. Specifically, Annex I enlists ‘robustness’ and ‘cybersecurity’ specifications for AI systems as standardisation deliverables to be developed. Once accepted by the European Commission, these technical standards become ’harmonized’ standards and compliance with the AIA will be presumed if providers adhere to them (Art. 40 AIA).
@@
The question arises as to why it is even necessary to regulate accuracy, robustness and cybersecurity of AI systems. One could assume that it is in the best interest of an economic actor to fulfill these requirements in the best possible way to gain a market advantage. The European Commission’s impact assessment of the AI Act recognizes this thought with respect to accuracy and robustness, stating that “an economic operator […] would anyway have to ensure that their product actually works” (European Commission, 2021, Annex 4). However, the impact assessment clarifies that “is important that these requirements are included in the regulatory framework so that substandard operators need to improve their procedures” (European Commission, 2021, Annex 4).
Therefore, the purpose of the specific requirements set out in Art. 15 AIA is to achieve the overarching objective of ensuring the trustworthiness of AI systems (see Art. 1(1) AIA) and to advance the cybersecurity agenda of the EU. [^7] COM/2010/245 final/2, COM/2021/118 final.
@@ 5. Requirements for High-Risk AI Systems
Our findings are supported by an historic analysis of the legislation process. As outlined in Section 4, the AIA builds on the Ethics Guidelines for Trustworthy AI (AI IHEG, 2019). In the guidelines, the principle of ’technical robustness and safety’ includes resilience against attacks, but does not mention cybersecurity. The White Paper on Artificial Intelligence (European Commission, 2020), which elaborates on these guidelines, still lists resilience to attacks against AI systems under “robustness and accuracy” without differentiating those terms from cybersecurity. The first official draft of the AIA by the European Commission [^10] COM/2021/206 final. was the first official document to distinguish between these three terms and assigned “resilience against attacks” to cybersecurity rather than robustness. Rec. (27), which refers to the IHGE guidelines, seems to be a remnant of this development process. It demands under the term ‘technical robustness’ that AI systems should be resilient “against attempts to alter the use or performance of the AI system”, essentially asking for adversarial robustness.
@@
The AIA regulates AI systems, but not AI models, with the only exception being GPAIMs. ML research, in contrast, often focuses on developing technical solutions for ML models. This raises the question of whether solely relying on technical solutions for ML models is enough to ensure the compliance of a HRAIS with Art. 15 AIA—or whether additional measures are needed. Rec. (97) specifies that an AI model is an essential component of an AI system. [^11] Although Rec. (97) specifically refers to GPAIMs, the wording suggests that the statement about the relationship between AI systems and AI models is of a general nature. Additional components can include, i.a., user interfaces, sensors, databases, network communication components, or pre- and post-processing mechanisms for model in- and outputs (Rec. (97), (Junklewitz et al., 2023)). All these individual components should contribute to the overall robustness of the AI system, particularly in scenarios where some components may fail. This is illustrated by Art. 15(4)(ii) AIA, which states that robustness may be ensured through technical redundancy solutions, including “back-up or contingency plans”. Furthermore, Art. 15(5)(iii) AIA stipulates that the cybersecurity of AI systems shall be achieved through technical solutions that, “where appropriate”, target training data, pre-trained components, the AI model or its inputs. This binding provision suggests that at least these different components of the AI system are required to be assessed individually for their appropriateness in mitigating cybersecurity attacks. Thus, Art. 15 AIA should not be understood as requiring a single, unified assessment of the requirements. Instead, it must be interpreted as mandating that each component, including one or more ML models, be assessed individually. The assessment of the AI system’s overall performance is then derived from an aggregation of the individual performance results (Kumar and Anderson, 2023). This requires an interdisciplinary approach that draws on expertise from fields such as ML, engineering, and human-computer interaction. To establish a common understanding, it can prove beneficial to formally describe the evaluation process of an entire AI system, including potential challenges, such as interdependencies of technical solutions.
@@
In ML, robustness is often measured using an accuracy metric. Typically, this involves comparing the accuracy (or error rates) evaluated on an unperturbed dataset from the original distribution with the accuracy on a perturbed test set (e.g., sampled from the shifted distribution or containing adversarial samples) (Taori et al., 2020; Hendrycks et al., 2021; Goodfellow et al., 2015). The smaller the difference between these two accuracy results, the better the robustness. The choice of the accuracy metric thus has an impact on the measurement of robustness. As a result, the ML model may appear more robust under some accuracy metrics than others. The selection of favorable metrics has been studied in fair ML under the term fairness hacking (Meding and Hagendorff, 2024; Simson et al., 2024; Black et al., 2024). Without entering into the debate, we note that there is an ongoing discussion in the ML literature about the existence and characteristics of a trade-off between robustness and accuracy. While some research showed that enhancing robustness leads to a drop in test accuracy (Zhang et al., 2019; Rade and Moosavi-Dezfooli, 2022; Tsipras et al., 2019), others believe that robustness and accuracy are not conflicting goals and can be achieved simultaneously (Yang et al., 2020; Raghunathan et al., 2020). Technical standards and guidelines by the EU Commission should provide instructions on how AI system providers should choose an appropriate ‘accuracy’ measure, especially when it is used to assess robustness in subsequent steps.
@@
Art. 15(4)(iii) AIA states that AI systems must be explicitly developed in such a way that they “duly address” feedback loops and “eliminate or reduce” the risks associated with them. According to Rec. (67), feedback loops occur when the output of an AI system influences its input in future operations, an under understanding that aligns with the concept as found in the ML literature. Feedback loops are a well-studied problem manifesting in various forms (Pagan et al., 2023), with the most common issues being a distribution shift (Perdomo et al., 2020) or a selection bias (Kilbertus et al., 2020; Lum and Isaac, 2016). Importantly, in this context, the risk of ’biased outputs’ in feedback loops (Art. 15(4)(iii) AIA) is often studied in the literature on fairness in ML rather than in the literature on robustness in ML, which traditionally constitute different research fields and communities (Lee et al., 2021).  [^14] For example, whether there is a trade-off between robustness and fairness, or if both pursue similar goals, remains an active discussion in the ML community (Lee et al., 2021; Xu et al., 2021; Pruksachatkun et al., 2021).
@@
Mandating a cybersecurity level that is ‘appropriate to the relevant circumstances’ acknowledges that complex ML models generally cannot be expected to be fully resistant to all types of adversarial attacks. This has two major reasons: First, it is impossible to anticipate all types of possible attacks. This is acknowledged by Art. 9(5) AIA which states that measures adopted under the risk management system are not expected to remove all existing risks. Second, complete protection against a specific attack cannot be guaranteed, especially as adversaries continuously adapt their strategies to overcome possible defense mechanisms (Xie et al., 2023; Kumar et al., 2023). The appropriateness of a certain performance level must consider the intended purpose of the system and the generally acknowledged SOTA (see Art. 8(1) AIA). The measures to ensure cybersecurity adopted are not expected to eliminate all existing risks, but the overall residual risk must be acceptable (see Art. 9(1) and (4) AIA). Thus, when determining the appropriateness of technical solutions, all applicable requirements of the AIA must be balanced, while also mitigating risks to health, safety, and fundamental rights.
@@
Art. 15(5) AIA differentiates between ’system vulnerabilities’ (Art. 15(5)(i) AIA) and ’AI-specific vulnerabilities’ (Art. 15(5)(iii) AIA). As the term vulnerability is not defined, we provide a working definition. The United States’ Common Vulnerabilities and Exposures (CVE) system defines vulnerability as “[a]n instance of one or more weaknesses […] that can be exploited, causing a negative impact to confidentiality, integrity, or availability” (CVE, 2024). Art. 15(5)(iii) AIA provides a non-exhaustive list of components of an AI system that expose AI-specific vulnerabilities, such as training data, pre-trained components used in training, inputs, or the AI model. However, there might be additional components of the AI system that may also harbor AI-specific vulnerabilities. The question is how to identify them. We suggest performing a hypothetical test. AI models play a central role in an AI system. If a vulnerability would be eliminated by replacing the AI model with a non-AI model, it should be deemed ‘AI-specific’. To define a non-AI model, we return to the definition of an AI system under the AIA. It has been argued that the central characteristic of an AI system is its ability to infer from input to output (Hacker, 2024). This inference ability is typically performed by one or more AI models within an AI system. Therefore, non-AI models are all models lacking inference capability, such as rule-based decision-making systems that rely on predefined rules and logic defined by human experts. [^16] Note that non-AI rule-based systems use human-defined rules, while rule-based ML models infer rules from data (Naik et al., 2023; Weiss and Indurkhya, 1995), qualifying as AI models. In a different context, the AI IHEG ethics guidelines (AI IHEG, 2019) suggest fallback plans where AI systems switch from a statistical (ML) approach to a rule-based or human-in-the-loop approach. Since AI-specific vulnerabilities relate to specific components of an AI system, we suggest viewing them as a subset of system vulnerabilities. To enhance clarity, technical standards should define both terms and mandate a process for identifying them.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732020isbn: 979-8-4007-1482-5/2025/06ccs: Social and professional topics Governmental regulationsccs: Security and privacy Software and application securityccs: Applied computing Law
@@ 1. Introduction
The European Union (EU) recently adopted the Artificial Intelligence Act (AIA) [^1] EU Regulation 2024/1689, 12.7.2025. which creates a legal framework for the development, deployment, and use of “human-centred and trustworthy artificial intelligence (AI)” (Art. 1 AIA). The AIA outlines desirable “ethical principles” of AI systems (Rec. (27)) and, i.a., imposes some of these as legally binding requirements for high-risk AI systems (HRAIS), e.g, AI systems intended to be used to take university admission decisions, or to evaluate individuals’ creditworthiness, and for general-purpose AI models (GPAIMs), e.g., multimodal large language models. While the AIA is recognized as being one of the first legally binding regulatory frameworks for AI (Chee and Hummel, 2024), it has faced criticism for its imprecise and incoherent terminology (Laux et al., 2024; Bomhard and Siglmüller, 2024), which will complicate its practical implementation. Previous work has examined the AIA and its legislative history to clarify terms like explainability (Bordt et al., 2022; Vitali, 2022; Pavlidis, 2024) and fairness (Deck et al., 2024). So far, little attention has been paid to robustness and cybersecurity. Only AI systems classified as high-risk (HRAIS) must meet the robustness and cybersecurity requirements set out in Art. 15 AIA. This paper thus focuses on requirements for HRAIS. To provide a clearer understanding, we compare these requirements with requirements for specific AI models, namely for GPAIMs with systemic risk, in Art. 55 AIA.
Technical solutions to ensure the robustness and cybersecurity of AI systems are often developed within the ML domain. Therefore, it is essential to inform ML research about the legal requirements to ensure compliance with the AIA. However, the vagueness of requirements for cybersecurity and robustness under the AIA makes it challenging to inform ML practitioners about the specific legal requirements to further the development of solutions that can ensure compliance with the AIA.
@@
A common understanding between technical and legal domains can be facilitated through technical standards. While the AIA sets out general rules, technical standards specify these rules in detail. Standards are technical specifications designed to provide voluntary technical or quality specifications for current or future products, processes or services. [^2] Art. 1, 2(1) EU Regulation 1025/2012, OJ L 316, 14.11.2012. They prescribe technical requirements, including characteristics such as quality or performance levels, terminology, and test methods. [^3] Art. 2(4)(a) and (c) ibid. Standards have long been integral to EU product legislation under the New Legislative Framework, upon which the AIA is built (Gorywoda, 2009). If approved by the EU Commission, technical standards become harmonized technical standards, which grants a presumption of conformity to products or processes that adhere to them. Consequently, compliance with these standards is deemed to fulfil the requirements of the AIA, thereby incentivising providers to adopt them (Art. 40 AIA). The development of harmonized technical standards for the AIA has been initiated by the EU Commission and is expected to be completed in the next few years. In addition, the EU Commission is also tasked with developing additional guidelines on the practical implementation on the application of Art. 15 AIA (Art. 96(1)(a) AIA).
@@
  • We analyse and explain the legal requirements related to robustness and cybersecurity in the AIA, identify related shortcomings, and offer possible solutions for some of these shortcomings.
@@
Our research employs a doctrinal analysis and as such involves a systematic examination and interpretation of provisions of the AIA, starting with Art. 15 AIA and extending to other provisions within the AIA. We refer to recitals in our analysis as they are likely a starting point of any future judicial engagement with these legal provisions. Indeed, whereas recitals are not strictly legally binding they have been highly authoritative in Court of Justice of the European Union (CJEU) case law. While our analysis is primarily grounded in legal scholarship, we map our findings to machine learning (ML) literature to bridge the gap between the legal and ML domains. This approach aims to inform a broader audience about the regulatory challenges and opportunities at the intersection of these fields. Our analysis follows the principle that law should be interpreted autonomously, recognizing that legal terminology does not necessarily align with technical terminology.
This paper is structured as follows: Section 2 provides a short background on robustness and cybersecurity in the ML literature. Section 3 provides an introduction to the AIA and Art. 15 AIA. Section 5 analyses the requirements outlined in Art. 15 AIA for HRAIS, addressing both general challenges pertinent to robustness and cybersecurity, as well as specific issues related to each requirement. Section 6 examines the requirements in Art. 55 AIA relevant to GPAIMs with systemic risk. Section 7 concludes with a summary and recommendations for future research.
@@ 2. An ML Perspective on Robustness and Cybersecurity
From a technical standpoint, adversarial robustness is one aspect of cybersecurity. Research on cybersecurity focuses on developing defences that protect computer systems from attacks compromising their confidentiality, integrity, or availability (Dasgupta et al., 2022). This encompasses aspects like data storage, information access and modification, and secure data transmission over networks (Sarker et al., 2021). Unlike robustness, cybersecurity is not a stand-alone concept in ML, but is discussed more broadly as both a tool for ensuring cybersecurity and a potential source of cybersecurity risks. [^4] E.g., NeurIPS’18 Workshop on Security in ML (Papernot et al., 2018), ICML’22 Workshop on ML for Cybersecurity (Applebaum et al., 2022). ML algorithms can be employed to detect and mitigate cybersecurity threats (Sarker et al., 2021), but can also introduce specific vulnerabilities that adversaries may exploit, such as data poisoning or adversarial attacks (Roshanaei et al., 2024; Rosenberg et al., 2021).
@@
[…]
[A flowchart illustrating how different types of inputs influence a machine learning (ML) model’s performance.]A test example sampled from the original population distribution is labelled as a safe example and, when passed to the ML model, results in a correct output. If this example is maliciously perturbed, it becomes an adversarial example, which causes the ML model to produce an incorrect output. Separately, a distribution shift on the population distribution leads to a shifted population distribution. Sampling from this shifted population distribution produces a non-adversarial example, which, when passed to the ML model, also leads to an incorrect prediction.
@@
[Overview of mapping between the Machine Learning (ML) domain and the Artificial Intelligence Act (AIA) domain.]The image shows two non-overlapping circles, one representing the Machine Learning (ML) domain and one the Artificial Intelligence Act (AIA) domain. The ML domain includes adversarial robustness and non-adversarial robustness. Arrows connect these concepts to the AIA domain. Adversarial robustness is linked to Cybersecurity, Art. 15(5) AIA. Non-adversarial robustness connects to Robustness, Ar. 15(4) AIA.
@@
The AIA creates harmonized rules for certain AIA systems in order to incentivize the use of such systems in the EU market and prevent regulatory fragmentation between member states (Rec. (1)). It is formally structured into recitals (Rec.), articles (Art.), and annexes. Recitals are not strictly legally binding and outline the rationale behind the articles, articles delineate specific binding obligations, and the annexes provide additional details and specifications to support the articles  (Klimas and Vaiciukaite, 2008). [^5] From now on, whenever we cite recitals, we refer to those in the AIA without explicitly indicating it. Art. 3(1) AIA defines an AI system as “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs […] that can influence physical or virtual environments”. These AI systems are regulated differently based on their perceived risk level (Bomhard and Siglmüller, 2024; Sioli, 2021): Those posing unacceptable risks, such as social scoring, are prohibited or subject to qualified prohibitions; HRAIS, such as those used in medical devices, are allowed but must comply with certain requirements and undergo pre-assessment; other AI systems are subject only to specific transparency and information obligations. Among these categories, only HRAIS must fulfil the robustness and cybersecurity requirements under Art. 15 AIA. According to Art. 16(a) AIA, providers of HRAIS must ensure compliance with these requirements. An AI system is considered a HRAIS if it is either a safety component of a product or a product itself regulated under specific legislation, such as medical devices, machinery, or toys (Art. 6(1) AIA, Annex I), or if it poses a significant risk of harm to the health, safety, or fundamental rights of individuals in specific areas, such as education, employment, or law enforcement (Art. 6(2) and (3) AIA, Annex III).
@@
In addition to AI systems, the AIA establishes a separate regime of legal requirements in chapter V of the AIA for a very specific type of AI models, namely GPAIM (e.g., multimodal large language models, see also Section 6). Note that the AIA differentiates between AI systems and AI models. While Rec. (97) specifies that an AI model is an essential component of an AI system, there is no clear legal delineation of the two concepts, as we discuss further in Section 5.1. GPAI models are AI models that can perform tasks that they were not originally trained for (Gutierrez et al., 2023), such as large language models (OpenAI, 2023; Gemini Team et al., 2023), or large text-to-image models (Ramesh et al., 2022). Other types of AI models are not regulated and mentioned in the AIA. These are models that are created with a specific objective and can only accomplish tasks they are trained to perform (e.g., translation, classification).
@@
The AIA does not directly define specific technical requirements. Instead, it sets out ‘essential requirements’ that AI systems must comply with, and which are concretized by so-called technical standards. The regulatory concept of relying on standardization is a well-established process in EU product legislation and is referred to as the New Legislative Framework (Gorywoda, 2009). It traditionally involves the participation of stakeholders, such as providers of AI systems. The European Commission has already issued a standardization request to the European Committee for Standardisation and the European Committee for Electrotechnical Standardisation to develop standards for the AIA until 30 April 2025.666C(2023)3215 - Standardisation request M/593. Specifically, Annex I enlists ‘robustness’ and ‘cybersecurity’ specifications for AI systems as standardisation deliverables to be developed. Once accepted by the European Commission, these technical standards become ‘harmonized’ standards and compliance with the AIA will be presumed if providers adhere to them (Art. 40 AIA).
@@
The impact assessment clarifies that “is important that these requirements are included in the regulatory framework so that substandard operators need to improve their procedures” (European Commission, 2021, Annex 4). Therefore, the purpose of the specific requirements set out in Art. 15 AIA is to achieve the overarching objective of ensuring the trustworthiness of AI systems (see Art. 1(1) AIA) and to advance the cybersecurity agenda of the EU. [^7] COM/2010/245 final/2, COM/2021/118 final.
In particular, Art. 15 AIA interacts with numerous other legal requirements that impose direct or indirect cybersecurity-related obligations on products that can qualify as HRAIS, such as the Machinery Regulation, the Medical Device Regulation, and the European Health Data Space. While cybersecurity has become a central concern for lawmakers in the EU, the arising frameworks create a patchwork of legislation that is difficult to disentangle. Art. 15 AIA adds another layer to cybersecurity-related obligations in the EU.
@@ 5. Requirements for High-Risk AI Systems
Our findings are supported by an historic analysis of the legislation process. As outlined in Section 4, the AIA builds on the Ethics Guidelines for Trustworthy AI (AI IHEG, 2019). In the guidelines, the principle of ‘technical robustness and safety’ includes resilience against attacks, but does not mention cybersecurity. The White Paper on Artificial Intelligence (European Commission, 2020), which elaborates on these guidelines, still lists resilience to attacks against AI systems under “robustness and accuracy” without differentiating those terms from cybersecurity. The first official draft of the AIA by the European Commission [^10] COM/2021/206 final. was the first official document to distinguish between these three terms and assigned “resilience against attacks” to cybersecurity rather than robustness. Rec. (27), which refers to the IHGE guidelines, seems to be a remnant of this development process. It demands under the term ‘technical robustness’ that AI systems should be resilient “against attempts to alter the use or performance of the AI system”, essentially asking for adversarial robustness.
@@
The AIA regulates AI systems, but not AI models, with the only exception being GPAIMs. Accordingly, the wording of Art. 15 AIA refers to “systems”. ML research, in contrast, often focuses on developing technical solutions for ML models. This raises the practical question of whether solely relying on technical solutions for ML models is enough to also ensure the compliance of a HRAIS with Art. 15 AIA—or whether and how cybersecurity and robustness should be evaluated at the system level.
Rec. (97) specifies that an AI model is an essential component of an AI system. [^11] Although Rec. (97) specifically refers to GPAIMs, the wording suggests that the statement about the relationship between AI systems and AI models is of a general nature. Additional components can include, i.a., user interfaces, sensors, databases, network communication components, or pre- and post-processing mechanisms for model in- and outputs (Rec. (97), (Junklewitz et al., 2023)). All these individual components should contribute to the overall robustness of the AI system, particularly in scenarios where some components may fail. This is illustrated by Art. 15(4)(ii) AIA, which states that robustness may be ensured through technical redundancy solutions, including “back-up or contingency plans”. Furthermore, Art. 15(5)(iii) AIA stipulates that the cybersecurity of AI systems shall be achieved through technical solutions that, “where appropriate”, target training data, pre-trained components, the AI model or its inputs. This binding provision suggests that at least these different components of the AI system are required to be assessed individually for their appropriateness in mitigating cybersecurity attacks. Thus, Art. 15 AIA should not be understood as requiring a single, unified assessment of the requirements. Instead, it must be interpreted as mandating that each component, including one or more ML models, be assessed individually. The assessment of the AI system’s overall performance is then derived from an aggregation of the individual performance results (Kumar and Anderson, 2023). This requires an interdisciplinary approach that draws on expertise from fields such as ML, engineering, and human-computer interaction. To establish a common understanding, it can prove beneficial to formally describe the evaluation process of an entire AI system, including potential challenges, such as interdependencies of technical solutions.
@@
In ML, robustness is often measured using an accuracy metric. Typically, this involves comparing the accuracy (or error rates) evaluated on an unperturbed dataset from the original distribution with the accuracy on a perturbed test set (e.g., sampled from the shifted distribution or containing adversarial samples) (Taori et al., 2020; Hendrycks et al., 2021; Goodfellow et al., 2015). The smaller the difference between these two accuracy results, the better the robustness. The choice of the accuracy metric thus has an impact on the measurement of robustness. As a result, the ML model may appear more robust under some accuracy metrics than others. The selection of favourable metrics has been studied in fair ML under the term fairness hacking (Meding and Hagendorff, 2024; Simson et al., 2024; Black et al., 2024). Without entering into the debate, we note that there is an ongoing discussion in the ML literature about the existence and characteristics of a trade-off between robustness and accuracy. While some research showed that enhancing robustness leads to a drop in test accuracy (Zhang et al., 2019; Rade and Moosavi-Dezfooli, 2022; Tsipras et al., 2019), others believe that robustness and accuracy are not conflicting goals and can be achieved simultaneously (Yang et al., 2020; Raghunathan et al., 2020). Technical standards and guidelines by the EU Commission should provide instructions on how AI system providers should choose an appropriate ‘accuracy’ measure, especially when it is used to assess robustness in subsequent steps.
@@
Art. 15(4)(iii) AIA states that AI systems must be explicitly developed in such a way that they “duly address” feedback loops and “eliminate or reduce” the risks associated with them. According to Rec. (67), feedback loops occur when the output of an AI system influences its input in future operations, an under understanding that aligns with the concept as found in the ML literature. Feedback loops are a well-studied problem manifesting in various forms (Pagan et al., 2023), with the most common issues being a distribution shift (Perdomo et al., 2020) or a selection bias (Kilbertus et al., 2020; Lum and Isaac, 2016). Importantly, in this context, the risk of ‘biased outputs’ in feedback loops (Art. 15(4)(iii) AIA) is often studied in the literature on fairness in ML rather than in the literature on robustness in ML, which traditionally constitute different research fields and communities (Lee et al., 2021).  [^14] For example, whether there is a trade-off between robustness and fairness, or if both pursue similar goals, remains an active discussion in the ML community (Lee et al., 2021; Xu et al., 2021; Pruksachatkun et al., 2021).
@@
Mandating a cybersecurity level that is ‘appropriate to the relevant circumstances’ acknowledges that complex ML models generally cannot be expected to be fully resistant to all types of adversarial attacks. This has two major reasons: First, it is impossible to anticipate all types of possible attacks. This is acknowledged by Art. 9(5) AIA which states that measures adopted under the risk management system are not expected to remove all existing risks. Second, complete protection against a specific attack cannot be guaranteed, especially as adversaries continuously adapt their strategies to overcome possible defence mechanisms (Xie et al., 2023; Kumar et al., 2024). The appropriateness of a certain performance level must consider the intended purpose of the system and the generally acknowledged SOTA (see Art. 8(1) AIA). The measures to ensure cybersecurity adopted are not expected to eliminate all existing risks, but the overall residual risk must be acceptable (see Art. 9(1) and (4) AIA). Thus, when determining the appropriateness of technical solutions, all applicable requirements of the AIA must be balanced, while also mitigating risks to health, safety, and fundamental rights.
@@
Art. 15(5) AIA differentiates between ’system vulnerabilities’ (Art. 15(5)(i) AIA) and ’AI-specific vulnerabilities’ (Art. 15(5)(iii) AIA). As the term vulnerability is not defined, we provide a working definition. The United States’ Common Vulnerabilities and Exposures (CVE) system defines vulnerability as “[a]n instance of one or more weaknesses […] that can be exploited, causing a negative impact to confidentiality, integrity, or availability” (CVE, 2024). Art. 15(5)(iii) AIA provides a non-exhaustive list of components of an AI system that expose AI-specific vulnerabilities, such as training data, pre-trained components used in training, inputs, or the AI model. However, there might be additional components of the AI system that may also harbour AI-specific vulnerabilities. The question is how to identify them. We suggest performing a hypothetical test. AI models play a central role in an AI system. If a vulnerability would be eliminated by replacing the AI model with a non-AI model, it should be deemed ‘AI-specific’. To define a non-AI model, we return to the definition of an AI system under the AIA. It has been argued that the central characteristic of an AI system is its ability to infer from input to output (Hacker, 2024). This inference ability is typically performed by one or more AI models within an AI system. Therefore, non-AI models are all models lacking inference capability, such as rule-based decision-making systems that rely on predefined rules and logic defined by human experts. [^16] Note that non-AI rule-based systems use human-defined rules, while rule-based ML models infer rules from data (Naik et al., 2023; Weiss and Indurkhya, 1995), qualifying as AI models. In a different context, the AI IHEG ethics guidelines (AI IHEG, 2019) suggest fallback plans where AI systems switch from a statistical (ML) approach to a rule-based or human-in-the-loop approach. Since AI-specific vulnerabilities relate to specific components of an AI system, we suggest viewing them as a subset of system vulnerabilities. To enhance clarity, technical standards should define both terms and mandate a process for identifying them.